### v2

# CompBench: Benchmarking Complex Instruction-guided Image Editing

Abstract
While real-world applications increasingly demand intricate scene manipulation, existing instruction-guided image editing benchmarks often oversimplify task complexity and lack comprehensive, fine-grained instructions. To bridge this gap, we introduce CompBench, a large-scale benchmark specifically designed for complex instruction-guided image editing. CompBench features challenging editing scenarios that incorporate fine-grained instruction following, spatial and contextual reasoning, thereby enabling comprehensive evaluation of image editing models’ precise manipulation capabilities. To construct CompBench, We propose an MLLM-human collaborative framework with tailored task pipelines. Furthermore, we propose an instruction decoupling strategy that disentangles editing intents into four key dimensions: location, appearance, dynamics, and objects, ensuring closer alignment between instructions and complex editing requirements. Extensive evaluations reveal that CompBench exposes fundamental limitations of current image editing models and provides critical insights for the development of next-generation instruction-guided image editing systems. The dataset, code, and models are available in comp-bench.github.io.
Figure 1: Examples of CompBench. The figure showcases diverse instruction-guided image editing tasks across nine categories: object addition, object removal, object replacement, multi-object editing, multi-turn editing, implicit reasoning, action editing, location editing and viewpoint editing).
1 Introduction
Recent advances in instruction-guided image editing have pursued user-friendly and efficient manipulation of visual content. While such systems aim to simplify complex editing workflows, real-world applications often demand intricate instructions including spatial relationships, appearance details, and implicit reasoning. This necessitates the development of models with comprehensive capabilities in visual grounding, contextual understanding, and complex reasoning, thereby presenting substantial challenges to existing methodologies. However, as demonstrated in Figure 2, existing instruction-guided image editing benchmarks, e.g., Emu Edit [29], MagicBrush [38], and ReasonEdit [14], exhibit critical limitations in assessing these essential capabilities, primarily in three aspects:
Lack of Scene Complexity. A key limitation of existing benchmarks is their insufficient scene complexity, which hampers the representation of the intricate visual structures inherent in real-world images. This limitation primarily stems from two factors: the prevalent use of synthetic images and the selective inclusion of oversimplified real images.
The first issue arises from the widespread use of text-to-image generation models, such as Stable Diffusion [27], in previous benchmark construction [41, 22]. Although convenient, the resulting synthetic scenes typically exhibit sparse spatial layouts, limited foreground object diversity, minimal occlusions, and overly simplified textures and lighting conditions.
Such artificial compositions lack the dense object interactions, natural clutter, and photorealistic qualities essential for evaluating practical editing capabilities.
Even when incorporating real images from datasets, such as COCO [19], current benchmarks often present oversimplified scenarios. Although more realistic than synthetic counterparts, these images frequently exhibit elementary compositions that are insufficient for evaluating models on complex spatial relationships and interactions among multiple objects.
This problem is further exacerbated by benchmark design choices, wherein creators often deliberately exclude highly complex scenes featuring heavy occlusions, intricate details, or dynamic elements due to the challenges they pose for ground truth construction. While this practice facilitates more controllable evaluation, it creates a concerning discrepancy between benchmark performance and real-world applicability. Consequently, image editing models may attain high metric scores on these relatively simplified benchmarks, yet remain inadequate for real-world editing tasks that demand advanced scene understanding and nuanced manipulation. For instance, in reasoning-based tasks, InstructPix2Pix [2] exhibits a notable performance decline on our CompBench compared with ReasonEdit [14], showing decreases of approximately 2.5 in PSNR, 0.02 in SSIM, and 0.4 in CLIP-Score.
Limited Instruction and Task Comprehensiveness. Beyond their oversimplified visual scenes, current benchmarks are further constrained by the narrow scope of editing instructions and tasks, failing to reflect the complexity of real-world user demands. Most existing datasets rely on simplistic, atomic-level instructions (e.g., “change the dog to a cat”) that lack contextual reasoning, and compositional logic typical of real user requests. In practice, editing instructions often involve: multi-object manipulations (“remove the dog on the left and the cat behind the chair”), spatial relationships (“add a man to the right front of the woman”), or dynamic scene reasoning (“make the man in white bend down more”). The absence of such complex instructions artificially inflates model’s performance in perception tasks, as systems are not rigorously tested on the full spectrum of challenges that may be encountered in real-world applications. Furthermore, existing benchmarks present the deprivation in task diversity. They rarely cover multi-object editing scenarios, where multiple entities must be manipulated simultaneously under specific relational or contextual constraints (e.g., “remove the man in white and the woman behind the man”). Those benchmarks also tend to ignore action editing tasks, where models are required to adjust dynamic states or actions rather than static attributes (e.g., “make the dog jump higher”). Without assessing these realistic and complex editing scenarios, they provide an incomplete and often misleading evaluation of model robustness and practical applicability.
Table 1: Comparison of existing image-editing datasets and benchmarks. Our benchmark supports seven core editing tasks, including multi-object editing, action editing and viewpoint editing, which are absent from most prior benchmarks. Scenario complexity is quantified by four indicators: Avg. Obj. (average number of objects per image), Avg. Cat. (average number of object categories per image), OCC (percentage of images that contain occluded objects), and OOF (percentage of images that contain out-of-frame objects). Across all four metrics, our benchmark exhibits the highest complexity, underscoring its suitability for rigorous evaluation.
Datasets / Benchmarks Size Types Task Complexity Local Multi-turn Multi-obj. Implicit Action Location Viewpoint Avg. Obj. Avg. Cat. Occ. Rate OOF Rate Datasets InstructPix2Pix [2] 313K 4 8.71 4.16 79.36 81.39 EditWorld [37] 8.6K 1 8.01 4.45 76.67 72.00 UltraEdit [45] 4M 9 7.68 4.70 75.30 78.10 SEED-Data-Edit [10] 3.7M 6 6.21 3.82 63.82 81.40 HQ-Edit [15] 197K 6 8.22 4.84 66.97 60.30 AnyEdit [41] 2.5M 25 6.95 4.37 60.45 57.20 Benchmarks MagicBrush [38] 10K 5 9.22 5.04 91.71 78.34 EMU_Edit [29] 8 8.38 5.19 78.51 83.60 Reason-Edit [14] 0.2K - 4.93 3.09 54.30 52.28 I2EBench [22] 2K 16 7.03 4.20 68.78 66.40 Ours* 3K 9 13.58 5.87 98.47 86.38
Figure 2: Comparison between current datasets or benchmarks and our CompBench. First row: failed cases of other benchmarks. These results fail to maintain background consistencies or introduce noticable artifacts into the editing region. Second row: Examples of other benchmarks. These cases lack scene complexity and instruction comprehensiveness. Third row: Examples of our CompBench. Our benchmark features complex real-world scenarios with precise instructions.
Deficiencies in Edited Image Quality. Another critical yet often overlooked limitation of current benchmarks lies in the suboptimal quality of their edited images. Many existing datasets exhibit two predominant issues that compromise their reliability: (1) instruction-alignment inaccuracies, where the edited output fails to precisely fulfill the specified modifications. (2) conspicuous visual artifacts, including geometric distortions, contextual inconsistencies in backgrounds, and semantic incoherencies in foreground objects. These quality deficiencies introduce substantial noise into performance evaluations, potentially leading to misleading assessments of model capabilities. Such artifacts raises concerns about whether current benchmarks can effectively discriminate between truly sophisticated editing systems and those that merely produce superficially plausible results.
To address the aforementioned issues, we introduce CompBench, the first large-scale benchmark for instruction-based image editing in complex scenarios, specific examples are illustrated in Figure 1. Our benchmark offers the following three major advantages:
Realistic and Complex Scene Composition. As shown in Table 1, Our benchmark encompasses scenes that embody the diverse and realistic complexities present in real-world settings. We compare CompBench with existing datasets and benchmarks, including InstructPix2Pix [2], MagicBrush [38], Emu-Edit [29], Reason-Edit [14], Edit-World [37], I2E-Bench [22], UltraEdit [45], SEED-Data-Edit [10], HQ-Edit [15], and AnyEdit [41]–across four dimensions: average number of objects, average number of object categories, overall object occlusion rate, and out-of-frame object rate. CompBench consistently surpasses prior benchmarks in all these metrics. Notably, our average number of objects per image is approximately 47.3% higher than the second best (MagicBrush [38]), demonstrating the heightened complexity and diversity of our scenes.
Comprehensive Task Coverage and High Difficulty Level. As depicted in Figure 4(a), CompBench encompasses five major categories, consisting of local editing, multi-editing, action editing, scene spatial editing, and complex reasoning, spanning a total of nine tasks. We provide detailed breakdowns of the six core capabilities these tasks require and analyze the degree to which our benchmark challenges each, which can be found in Appendix. Additionally, we propose an Instruction Decomposition Strategy to improve the clarity and precision of image editing instructions. Specifically, we introduce a structured framework that organizes editing instructions along four complementary dimensions: spatial positioning (e.g., “left of the table”), visual attributes (such as color or texture), motion states (e.g., “flying”), and object entities. This structured approach converts potentially ambiguous requirements into well-defined specifications without sacrificing the natural expressiveness of instructions. By systematically covering each aspect of an editing operation while preserving the flexibility of natural language, our framework produces instructions that are both intuitively understandable and technically precise for complex image editing tasks.
High-Quality Data Curation. Every sample in CompBench is meticulously constructed through multiple rounds of expert review, ensuring the highest quality of edits. Unlike other benchmarks where editing failures are common, all data in CompBench represent successfully executed editing results, with SSIM (Structural Similarity Index Measure) scores significantly outperforming those of other datasets, as illustrated in Figure 4(b). This rigorous quality control ensures that CompBench provides a reliable assessment of model performance in realistically complex editing scenarios.
2 Related Work
Instruction-guided Image Editing. Instruction-guided image editing enables efficient image manipulation using only textual editing instructions, without the need for manual mask drawing or providing explicit visual elements for editing. This approach aligns more closely with user intentions. Diffusion Models [11], particularly Stable Diffusion [27] (SD), facilitate this task significantly by supporting explicit text inputs. The advent of methods such as InstructPix2Pix [2], which are based on Diffusion Models [11], has greatly improved the effectiveness of instruction-based image editing. InstructPix2Pix [2] leverages large language models (LLMs) [32, 4, 3, 31] and text-to-image (T2I) [26, 25, 28, 27] models to generate large-scale datasets and trains a diffusion model that is capable of following natural language instructions. HIVE [44] introduces a reward model that leverages human feedback to align image editing outputs with human preferences. Approaches such as SmartEdit [14], MGIE [9], and the latest Step1X-Edit [21] interactively integrate image and instruction representations using multi-modal large language models (MLLMs) [18, 1, 20, 34], injecting these capabilities into diffusion models to better understand complex instructions and achieve more precise edits. AnyEdit [41] constructs an extremely large-scale multi-task dataset and adopts a mixture-of-experts (MoE) [8, 6] architecture to better accommodate diverse editing tasks. SEED-X [10] utilizes a visual tokenizer to unify image comprehension and generation, establishing a unified multi-granularity comprehension and generation model that enhances editing performance. GoT [7] employs Generation Chain-of-Thought [36] reasoning, incorporating stepwise reasoning into visual generation and editing processes, and allows users to explicitly follow the editing reasoning process, thereby enabling more refined edits.
Image Editing Benchmarks.  High-quality image editing datasets and benchmarks are crucial for both model training and evaluation. MagicBrush [38] manually constructs a 10K-scale instruction-based image editing dataset, encompassing single-turn, multi-turn, mask-provided, and mask-free editing tasks. EMU-edit [29] introduces a challenging benchmark comprising seven diverse editing tasks. HQ-Edit [15] employs a scalable data collection pipeline to create a high-quality dataset of 200K instruction-guided image editing samples. SmartEdit [14] introduces Reason-Edit, a small-scale, manually curated benchmark focused on complex instruction-based image editing. Edit-world [37] presents the concept of world-instructed image editing and creates a dataset featuring instructions in a world context. I2EBench [22] proposes a comprehensive benchmark for instruction-based image editing evaluation, including automated multi-dimensional assessment. UltraEdit [45] develops a scalable framework for producing large and high-quality image editing datasets, introducing a large-scale instruction-based dataset. SEED-Data-Edit [10] provides a hybrid dataset composed of auto-generated, real-world, and human-annotated multi-turn editing samples, demonstrating its effectiveness in advancing instruction-based image editing.
Figure 3: The construction pipeline of CompBench. The pipeline consists of two main stages: (a) Source data collection and preprocessing, wherein high-quality data are identified through image quality filtering, mask decomposition, occlusion and continuity evaluation, followed by thorough human verification. (b) Task-specific data generation using four specialized pipelines within our MLLM-Human Collaborative Framework, where multimodal large language models generate initial editing instructions that are subsequently validated by humans to ensure high-fidelity, semantically aligned instruction-image pairs for complex editing tasks.
3 CompBench
3.1 Task Categorization and Definitions
Our complex instruction-guided image editing benchmark, CompBench, contains 3k+ image-instruction pairs. To enhance the comprehensiveness of evaluation, we categorize editing tasks into five major classes with nine specific tasks based on their characteristics:(1) Local Editing: focuses on manipulating local objects, including object removal, object addition and object replacement. (2) Multi-editing: addresses interactions among multiple objects or editing steps, including multi-turn editing and multi-object editing. (3) Action Editing: modifies the dynamic states or interactions of objects. (4) Scene Spatial Editing: alters scene spatial properties, consisting of location editing and viewpoint editing. (5) Complex Reasoning: requires implicit logical reasoning, including implicit reasoning. Examples of these tasks can be found in Figure 1.
3.2 Dataset Generation
In this section, we detailedly demonstrate the generation process of our CompBench. The overall pipeline is shown in Figure 3.
Source Data Collection and Preprocessing. To address the scarcity of high-quality paired complex editing data, we select MOSE [5], a video instance segmentation (VOS) dataset featuring complex scenes with multi-object masks. The dataset undergoes a rigorous preprocessing pipeline: We first filter low-quality video frames using a mixture of no-reference image quality assessment metrics (e.g., NIQE [42]) to eliminate blurry, low-contrast, or corrupted samples. Then, A professional team manually verifies the filtered data, retaining only high-quality images. For mask preprocessing, multi-object masks are decomposed into single-object masks to isolate editable entities. A multimodal large language model (MLLM, e.g., Qwen-VL [34]) evaluates mask continuity and occlusion levels, discarding discontinuous and incomplete masks with high occlusion rate. Similarly, annotators further check these masks to ensure pixel-level precision.
Task-specific Data Generation Pipelines. To address the unique challenges and diversity of complex instruction-guided image editing tasks, we design four specialized data construction pipelines tailored to distinct task categories: (1) local editing pipeline for object-level manipulations (object removal, object addition, object replacement). (2) action/scene spatial editing pipeline for modifying object dynamics or scene perspectives (action editing, location editing, viewpoint editing). (3) complex reasoning pipeline for implicit contextual edits requiring reasoning (implicit reasoning). (4) multi-editing pipeline for multi-object and multi-turn editing tasks. All pipelines adopt a unified MLLM-Human Collaborative Framework: multimodal large language models (MLLMs) [18, 1, 20, 34] generate initial task-specific instructions by analyzing visual scenes and editing goals, followed by human validation to ensure instruction-image semantic alignment and image editing fidelity. Unsuccessful edits are iteratively re-generated or discarded, retaining only high-fidelity samples that satisfy both linguistic precision and visual realism. During inference, models directly process free-form natural language instructions without pipeline-specific constraints, eliminating intermediate processing steps and enhancing operational efficiency. Detailed implementation procedures for each pipeline are provided in the appendix A.
Figure 4: Characteristics and statistics of CompBench. (a) Task taxonomy of CompBench, illustrating the full range of task types. (b) SSIM [35] comparison among different datasets and benchmarks. Note that UltraEdit [45] and InstructPix2pix [2] are datasets, whereas the remaining entries are benchmarks.
Instruction Decomposition Strategy. To improve the construction of image editing instructions with greater clarity and precision, we propose a structured framework that organizes editing instructions along four complementary aspects: spatial positioning, visual attributes, motion states, and object entities. This structured approach transforms potentially ambiguous editing requirements into well-defined specifications while maintaining natural expressiveness. The method employs a two-phase generation process: first, multimodal large language models produce dimension-aware instruction candidates by analyzing visual contexts. Then human experts refine these to ensure precision and consistency. By systematically addressing each aspect of the editing operation while preserving the flexibility of natural language, this framework enables the creation of instructions that are both intuitively understandable and technically precise for complex image editing tasks.
Characteristics and Statistics. As illustrated in Figure 4(a), our benchmark comprises 5 major categories encompassing a total of 9 complex editing tasks, yielding 3k+ image editing samples with corresponding complex instructions. Since we lack global captions for pre-edited and post-edited images, we employ the Structural Similarity Index Measure (SSIM) [35] to evaluate the semantic consistency between image pairs as a quality assessment metric. As shown in Figure 4(b), CompBench achieves notably higher SSIM than other datasets and benchmarks.
Notably, our dataset features significantly more challenging editing tasks, each requiring comprehensive capabilities such as visual grounding and complex reasoning. Detailed analysis of core competencies essential for our benchmark are discussed in Appendix B. To systematically evaluate scene complexity, we adopt multiple quantitative indicators such as average number of scene objects and categories. These metrics demonstrate that our benchmark exhibits substantially higher complexity compared to existing benchmarks.
4 Experiments
4.1 Settings
Baselines. Given that our study specifically targets instruction-guided image editing tasks, we restrict our selection to instruction-guided image editing models and exclude approaches based on global description guidance. The evaluated models include: InstructPix2pix [2], MagicBrush [38], HIVE [44], Smart-edit [14], MGIE [9], HQ-Edit [15], CosXL-Edit [30], UltraEdit [45], AnyEdit [41], Seed-X-Edit [10], GoT [7], and the recently released Step1X-Edit [21].
Table 2: Evaluation results on local editing, multi-object editing and implicit reasoning. LC-T denotes local CLIP scores between the edited foreground and the local description. LC-I refers to the CLIP image similarity between the foreground edited result and ground truth (GT) image. Top-three evaluation results are highlighted in red (1st), blue(2nd), and green (3rd).
Model Local Editing Multi-object Editing Implicit Reasoning Foreground Background Foreground Background Foreground Background LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ InstructPix2pix [2] 19.445 0.777 21.416 0.695 0.137 19.624 0.779 20.200 0.658 0.162 19.007 0.793 21.806 0.683 0.125 MagicBrush [38] 20.067 0.798 23.415 0.744 0.088 19.878 0.800 23.524 0.727 0.092 19.526 0.828 22.145 0.714 0.106 HIVE-w [44] 19.820 0.770 19.908 0.641 0.198 20.024 0.775 19.594 0.609 0.224 18.634 0.777 20.268 0.602 0.219 HIVE-c [44] 19.236 0.772 21.741 0.689 0.147 19.585 0.781 21.560 0.663 0.155 18.938 0.786 22.168 0.666 0.132 Smart-edit-7B [14] 20.034 0.798 24.398 0.761 0.073 19.888 0.803 23.959 0.731 0.090 19.743 0.832 23.058 0.732 0.096 MGIE [9] 18.957 0.780 20.572 0.708 0.175 19.830 0.788 18.296 0.692 0.274 17.728 0.801 24.432 0.780 0.088 CosXL-Edit [30] 19.029 0.778 20.442 0.706 0.156 19.550 0.788 20.382 0.682 0.171 18.269 0.794 20.984 0.681 0.161 HQ-Edit [15] 18.316 0.734 12.240 0.419 0.441 19.163 0.757 12.987 0.412 0.421 18.864 0.767 12.321 0.396 0.452 UltraEdit [45] 19.618 0.786 22.938 0.783 0.145 20.022 0.795 22.326 0.719 0.164 18.350 0.784 23.374 0.717 0.145 AnyEdit [41] 19.932 0.794 22.769 0.714 0.125 19.875 0.809 22.789 0.697 0.129 19.588 0.816 20.271 0.639 0.191 SEED-X [10] 17.933 0.780 21.466 0.805 0.139 19.092 0.795 20.638 0.788 0.158 17.467 0.784 21.506 0.709 0.134 GoT [7] 20.268 0.807 24.675 0.890 0.067 19.919 0.804 21.296 0.826 0.127 19.237 0.820 24.738 0.860 0.088 Step1X-Edit [21] 20.501 0.817 23.371 0.882 0.078 20.213 0.828 22.696 0.873 0.089 19.312 0.850 23.435 0.869 0.082
Table 3: Evaluation results on multi-turn editing.
Model Turn1 Turn2 Foreground Background Foreground Background LC-T LC-I PSNR SSIM LPIPS LC-T LC-I PSNR SSIM LPIPS InstructPix2pix [2] 19.424 0.784 21.073 0.676 0.142 19.818 0.776 17.607 0.568 0.238 MagicBrush [38] 19.977 0.812 24.020 0.730 0.089 20.253 0.811 21.244 0.682 0.134 HIVE-w [44] 19.784 0.781 20.040 0.621 0.196 20.129 0.761 17.291 0.532 0.272 HIVE-c [44] 19.756 0.787 21.330 0.660 0.155 19.812 0.778 18.346 0.590 0.217 Smart-edit-7B [14] 19.876 0.817 24.632 0.740 0.080 20.050 0.807 23.404 0.724 0.104 MGIE [9] 19.355 0.801 21.563 0.731 0.143 19.695 0.798 18.382 0.655 0.223 HQ-Edit [15] 18.987 0.755 12.950 0.410 0.422 18.935 0.740 12.032 0.383 0.499 CosXL-Edit [30] 19.389 0.787 20.233 0.679 0.171 19.394 0.771 16.752 0.590 0.301 UltraEdit [45] 19.990 0.792 23.763 0.715 0.116 20.177 0.779 22.917 0.715 0.139 AnyEdit [41] 19.953 0.812 23.412 0.711 0.113 20.093 0.803 20.010 0.633 0.188 SEED-X [10] 19.139 0.795 21.042 0.792 0.153 18.729 0.753 13.793 0.457 0.404 GoT [7] 20.108 0.816 25.089 0.894 0.066 19.939 0.804 21.397 0.825 0.131 Step1X-Edit [21] 20.157 0.832 23.987 0.883 0.078 20.262 0.835 20.710 0.822 0.128
Evaluation Metrics and Methods. Evaluation metrics for image editing tasks must be well-suited to the complexity of our scenarios, providing a comprehensive and accurate assessment of editing performance in complex scenes. Moreover, the metrics should be tailored to reflect the unique characteristics of different task types. For tasks including local editing, multi-editing and implicit reasoning, we posit that an effective editing model should modify foreground objects while preserving background consistency. Therefore, we evaluate the editing performance from both foreground and background perspectives. For background consistency assessment, we compute three metrics, including PSNR, SSIM [35], and LPIPS [43], on the background regions. For the foreground evaluation, we consider two aspects: editing accuracy and instruction following. For editing accuracy, we measure the similarity between the edited result and the ground truth (GT) image by comparing their CLIP [24] image embeddings in the foreground region, thereby determining whether the edited foreground visually aligns with the GT. To assess instruction-following capability, we measure the CLIP [24] similarity between the edited foreground object and the textual description of the target region to evaluate the model’s ability to interpret and execute the given instructions. Additionally, for action editing, location editing and viewpoint editing tasks, where the object’s morphology, position, or viewpoint may change significantly, the aforementioned metrics are insufficient for evaluation. Thus, we employ GPT-4o [23] for assessment. Specifically, we design tailored prompts for these three tasks, instructing GPT-4o to evaluate the editing performance from multiple perspectives and assign a score between 0 and 10. The detailed prompt designs are provided in the Appendix D.
Table 4: Evaluation results on action editing, location editing and viewpoint editing tasks.
Model Action Location Viewpoint Average InstructPix2pix [2] 3.047 3.425 0.088 2.187 MagicBrush [38] 3.511 4.603 0.500 2.871 HIVE-w [44] 3.151 4.110 0.559 2.607 HIVE-c [44] 3.977 4.192 0.559 2.909 Smart-edit-7B [14] 4.233 3.890 1.412 3.178 MGIE [9] 1.921 1.726 0.059 1.235 HQ-Edit [15] 1.449 1.425 0 0.958 CosXL-Edit [30] 4.270 5.479 0.235 3.328 UltraEdit [45] 4.449 4.014 0.235 2.899 AnyEdit [41] 3.750 5.068 0.471 3.096 SEED-X [10] 2.270 3.028 1.412 2.237 GoT [7] 3.337 3.625 0.206 2.389 Step1X-Edit [21] 6.270 5.041 1.500 4.270
4.2 Experiment Results
The experimental results for local editing, multi-turn editing, multi-object editing, implicit reasoning, and action/location/viewpoint editing are presented in Tables 2,  3, and  4, respectively. Our key analysis of the results are as follows: (1) No model dominates across all tasks. Among all evaluated models, Step1X-Edit [21] emerges as the most prominent one, achieving top results in 17 out of 29 metrics (nearly 60%) across 9 tasks. Notably, Step1X-Edit [21], SmartEdit [14], and GoT [7] consistently deliver superior performance, securing top-three rankings in the majority of metrics across most tasks. In contrast, HQ-Edit [15] demonstrates substantially inferior results in nearly all tasks. (2) For multi-turn editing tasks, all models exhibit a notable decline in background consistency metrics during the second editing round. Among them, SmartEdit [14] maintains relatively robust performance in second editing turn. In multi-object editing, Step1X-Edit [21] outperforms other models by a significant margin, showcasing exceptional capability in simultaneous multi-object manipulation. (3) Beyond the consistently high-performing models, MGIE [9] also demonstrates competitive results in background consistency for implicit reasoning tasks. (4) For action/location/viewpoint editing tasks, GoT [7] does not perform as strongly compared to its results in other tasks. Step1X-Edit [21] remains the most reliable performer, achieving the only score above 6 (out of 10) in the highly challenging action editing task. Meanwhile, Smartedit [14] and CosXL-Edit [30] exhibit strong performance in these tasks, distinguishing themselves from other models.
5 Insights
In this section, we investigate the underlying factors that lead to varying performances among different models on our proposed CompBench, and offer perspectives on future research directions for the field of image editing.
Figure 5: Overall Model Performance.(a) Top 5 model performace in five major evaluation tasks. (b) Overall model performace across all tasks.
The Critical Role of MLLMs. Through systematic evaluation, we discover a strong correlation between architectural design and editing performance: multi-modal large language models (MLLMs) [18, 1, 20, 34, 13, 12, 40, 17] serve as a cornerstone for recent advances in this field. Specifically, for a fair comparison, we normalize and average the performance scores of all models across five major tasks, highlighting the top-5 models in Figure 5(a). Furthermore, we present the overall normalized results in Figure 5(b), aggregated across all tasks. The results reveal that Step1X-Edit [21] significantly outperforms others on complex instruction-editing tasks, closely followed by SmartEdit [14] and GoT [7]. Notably, while MGIE [9] ranks 10th out of 13 models overall, it demonstrates comparatively performance on implicit reasoning tasks (particularly background evaluation metrics) which require advanced reasoning capabilities. Interestingly, all these models, including Step1X-Edit, SmartEdit, GoT and MGIE, integrate MLLMs with diffusion models to better fuse textual and visual information. This integration enables these models to more accurately understand complex instructions and the overall visual context, ultimately driving superior editing performance in complex editing tasks.
The Importance of Reasoning Ability. In addition to incorporating MLLMs to enhance visual-text fusion, we have also recognized the significant impact of reasoning ability on editing performance. SmartEdit is trained on the reasoning segmentation dataset from LISA, which significantly enhances its reasoning capabilities and leads to outstanding results on multiple tasks. On the other hand, with the recent development of reasoning MLLMs, GoT introduces Chain-of-Thought (CoT) [36] into the editing process by leveraging MLLMs to generate reasoning chains. This approach further improves the model’s understanding of editing instructions and images, enabling more precise editing.
Our analysis reveals two critical insights that advance the field of complex instruction-guided image editing. Fisrt, MLLMs are indispensable for high-performance editing, as they uniquely bridge the gap between textual instructions and visual understanding. Second, multi-modal reasoning capabilities are foundational to achieving high-quality editing outcomes. These findings indicate that future research should focus on: (1) developing MLLM architectures specifically optimized for editing workflows rather than general vision-language tasks (2) advancing reasoning-aware training frameworks, such as reinforcement learning (RL)-based optimization of reasoning chains or utilizing reasoning data, to enhance editing precision and adaptability.
6 Conclusion
In this work, we introduce CompBench, the first large-scale benchmark specifically designed for comprehensive evaluation of instruction-guided image editing capabilities. Our meticulously constructed benchmark encompasses five major categories with nine specialized tasks targeting complex image manipulation scenarios, comprising over 3,000 high-quality image editing pairs with corresponding natural language instructions. We conduct extensive experimental evaluation across 13 state-of-the-art instruction-guided image editing models on all benchmark tasks to systematically assess the capabilities and limitations of contemporary editing systems and validate the efficacy of our evaluation framework. The experimental findings from CompBench not only reveal significant performance gaps in current models but also yield valuable insights that elucidate promising research directions for advancing next-generation image editing systems with enhanced reasoning abilities and fine-grained control.
Appendix A Implementation Details
A.1 Source Data Collection and Preprocessing
Figure 6: Prompts of Editing Evaluation.
The primary data collection and preprocessing pipeline has been described in section 3.2. For initial quality assessment of MOSE data, we employed four no-reference metrics (NIQE [42], MANIQA [39], MUSIQ [16], and CLIPIQA [33]). All images were systematically evaluated using these metrics, with the scores subsequently normalized and equally weighted to compute a composite value for filtering purposes. In terms of mask preprocessing, we decomposed multi-object masks into discrete single-object representations and utilized a multimodal large language model (specifically Qwen-VL [34]) to quantitatively assess mask continuity and occlusion levels, with the corresponding prompts illustrated in Figure 6. The comprehensive data preprocessing workflow underwent rigorous multi-round review and verification by a dedicated team of domain experts, thereby ensuring the exceptional quality and reliability of the final dataset.
A.2 Task-specific Data Generation Pipelines
Figure 7: Local Editing Pipeline.
Figure 8: Comparison of Different Blending Operations.
Figure 9: Action/Scene Spatial Editing Pipeline.
Figure 10: Complex Reasoning Pipeline.
Figure 11: Multi-editing Pipeline.
Due to the distinct characteristics of different tasks, we have designed specialized pipelines for data generation tailored to related task categories. The specific workflows for each pipeline are as follows:
Local Editing Pipeline for Object-Level Manipulations. As illustrated in Figure 7, given the preprocessed high-quality VOS dataset, we employ an inpainting model (PowerPaint [46]) to execute object removal based on precise object masks. The resultant outputs undergo rigorous evaluation and refinement through a Multimodal Large Language Model (MLLM) [18, 1, 20, 34] in conjunction with manual verification. For instruction generation, we provide the pre- and post-edited images alongside the corresponding masks, utilizing the MLLM to generate preliminary instructions, which are subsequently refined manually to ensure they accurately reflect the specific editing operations. To facilitate comprehensive evaluation, we systematically select and manually annotate captions for the mask regions before and after editing, thereby generating the definitive Object Removal dataset. For Object Addition data, we strategically reverse a subset of the Object Removal data and similarly employ MLLM complemented by manual annotations to construct the corresponding instructions. For Object Replacement, we implement a candidate target object generation approach, wherein the MLLM analyzes the before-and-after images and masks from Object Removal to propose plausible, diverse replacement objects that maintain contextual coherence with the scene’s characteristics. Following manual selection of appropriate objects, we execute object replacement using the inpainting model, with instructions formulated analogously to those in the aforementioned tasks.
To further enhance background consistency and overall quality of inpainting results, we implement a composite post-processing strategy that integrates dilation, Gaussian blurring, and Poisson blending. Specifically, we first dilate the edges of the target object in the post-editing image, employing a kernel size of 20. Subsequently, Gaussian blurring is applied to the dilated edges with a kernel size of 15 and a $\sigma_{X}$ value of 3. Finally, Poisson blending is applied between the original image and the modified object region in the post-editing image to achieve the definitive result. As illustrated in Figure 8, the unprocessed edited images (direct inpainting outputs) exhibit noticeable blurriness and diminished clarity. Initially, we explored a rudimentary blending approach, directly merging the edited region with the background from the original image. However, this methodology resulted in pronounced and conspicuous boundaries, compromising the overall blending quality. By incorporating dilation and Gaussian blurring, the boundary integration improved substantially, yet unnatural chromatic discrepancies between the edited region and the surrounding background persisted. To address this limitation, we further integrated Poisson blending, which generated significantly superior results, yielding natural and seamless integration without perceptible artifacts.
Action/Scene Spatial Editing Pipeline. For action editing and scene spatial editing tasks (including location and viewpoint editing), as illustrated in Figure 9, we strategically select pertinent data from the VOS dataset. Specifically, for action editing, we extract frames from the same video sequence where the background remains consistent, while the objects undergo motion transformations. For location editing, we identify frames from identical video sequences where the background maintains constancy, but the object exhibits positional displacement. For viewpoint editing, we select frames exhibiting perspectival variations within the video. The generation of editing instructions for these tasks follows a methodology analogous to the local editing pipeline, integrating MLLM capabilities with meticulous manual refinement.
Complex Reasoning Pipeline. As depicted in Figure 10, for complex reasoning data, we systematically select a subset of data from the original VOS dataset that demonstrates suitability for implicit contextual edit instructions, subsequently modifying these instructions through manual intervention to produce the definitive dataset.
Multi-Editing Pipeline. For multi-object/multi-turn editing illustrated in Figure 11, we composite results from the object removal subset within the VOS dataset, where multiple object masks coexist within the same video frame, to generate comprehensive multi-object removal outputs. Conversely, this approach establishes the foundation for multi-object addition. For multi-turn editing, the outcome from manipulating a single object constitutes an initial editing iteration, while the multi-object composite result represents a subsequent editing phase, culminating in the multi-turn removal outcome. Similarly, this procedural framework can be applied to multi-turn addition operations. All associated instructions undergo rigorous manual development and refinement to ensure precision and clarity.
A.3 Experiment Setting Details
Table 5: Configurations Details of Editing Models.
Method Configuration
InstructPix2Pix [2] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
MagicBrush [38] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
[…]
Smart-edit [14] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
MGIE [9] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
HQ-Edit [15] guidance_scale=7, num_inference_steps=30 image_guidance_scale=1.5
CosXL-Edit [30] guidance_scale=7.5, num_inference_steps=20 image_guidance_scale=1.5
UltraEdit [45] guidance_scale=7.5, num_inference_steps=50 image_guidance_scale=1.5
AnyEdit [41] guidance_scale=3, num_inference_steps=100 image_guidance_scale=3
SEED-X [10] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
GoT [7] guidance_scale=5.0, num_inference_steps=50 image_guidance_scale=1.0
Step1X-Edit [21] guidance_scale=6, num_inference_steps=28
Evaluation Configurations. We basically use the official settings of all models during evaluation. Specific configurations of each model are shown in Table 5.
GPU Usage. All experiments using the inpainting model to construct our benchmark were conducted on 8 NVIDIA A6000 GPUs (48 GB each). The same setup was used during evaluation to meet the memory and computation demands of all editing models.
Appendix B Capability Analysis of CompBench
Table 6: Task Competence Analysis. We identify six core competencies essential for complex image editing. The degree to which each competency is required by a given task is classified as Low (L), Medium (M), High (H), or Ultra High (UH, bolded for extraordinarily high requirements).
Task Visual Appearance Relation Complex Scene 3D Grounding Control Understanding Reasoning Consistency Geometry Object Removal H H H H UH L Object Addition H UH H H UH L Object Replacement H UH H H UH L Action Change M H H M UH H Location Change H M M H UH M Viewpoint Change M M M M UH UH Multi-turn Editing H UH H H UH M Multi-object Editing UH UH UH H UH H Implicit Reasoning H H UH UH UH H
To comprehensively characterize the requirements of complex image editing, we identify six core competencies that a sophisticated editing system must demonstrate: (1) Visual Grounding. the precise localization of target objects or regions; (2) Appearance Control. fine-grained manipulation of visual attributes such as color, texture, and illumination; (3) Relation Understanding. accurate modeling of semantic and spatial dependencies among objects; (4) Complex Reasoning. implicit logical deduction from contextual cues; (5) Scene Consistency. holistic preservation of spatial layout, occlusion patterns, and contextual coherence; and (6) 3D Geometry. understanding and manipulating three-dimensional structure and viewpoint.
The quantitative correspondence between individual tasks and the competencies they necessitate is summarized in Table 6. Local editing tasks place an exceptionally high premium on Scene Consistency, as seamless integration of the modified region is paramount. Multi-object Editing requires a balanced proficiency in Appearance Control, Relation Understanding, and Scene Consistency to effectively manage complex inter-object interactions. Viewpoint Editing uniquely depends on the 3D Geometry competency to facilitate perspective transformations, whereas Implicit Reasoning imposes stringent demands on Complex Reasoning to infer indirect editing intentions.
Overall, CompBench presents substantial challenges by requiring the simultaneous integration of these multi-dimensional competencies, thereby reflecting the intricacy of real-world image-editing scenarios.
Appendix C Cases of CompBench
Figure 12: Cases of Local Editing in CompBench.
Figure 13: Cases of Multi-editing in CompBench.
Figure 14: Cases of Action, Location and Viewpoint Editing in CompBench.
Figure 15: Cases of Local Editing Results.
Figure 16: More Cases of Local Editing Results.
Figure 17: Action, Location and Viewpoint Editing Examples.
Figure 18: Action, Location and Viewpoint Editing Examples.
Figure 19: Action, Location and Viewpoint Editing Examples.
In this section, we present additional exemplars from CompBench and comprehensive evaluation results. In Figure 12, we demonstrate representative instances of local editing operations (object addition, object removal, and object replacement) within our CompBench. In Figure 13, we illustrate selected cases of multi-turn and multi-object editing outcomes in our benchmark. In Figure 14, we showcase exemplary instances of action editing, location editing, and viewpoint editing capabilities within our benchmark.
We further present qualitative results from all evaluated models on our benchmarks. The comparative editing outputs across all models for local editing, multi-editing, and implicit reasoning tasks can be examined in Figures 15 and 16. The corresponding results for action editing, location editing, and viewpoint editing are displayed in Figures 17,18, and19. Detailed evaluation protocols and analytical discussions are presented in Section D.
Appendix D Evaluation Details
In this section, we delineate the comprehensive evaluation methodology and procedural framework employed in our assessment.
For tasks encompassing local editing, multi-editing, and implicit reasoning, we require models to modify foreground elements while maintaining background fidelity. Consequently, we evaluate editing performance from both foreground and background perspectives. Background consistency is quantitatively assessed utilizing PSNR, SSIM, and LPIPS metrics on background regions, while foreground evaluation incorporates both editing accuracy via CLIP image embedding similarity with ground truth exemplars and instruction adherence via CLIP Score with the local foreground caption. The foreground captions are meticulously annotated and validated, with representative examples illustrated in Figure 12.
To objectively quantify the quality of Action Editing, Location Editing, and Viewpoint Editing, which require substantial image manipulations, we devise an automated evaluation pipeline leveraging GPT-4o: for each source–edited image pair, the pipeline generates a reproducible and interpretable composite score ranging from 0 to 10, accompanied by a concise textual justification. Initially, we formulate three task-specific prompts as demonstrated in Figure 6, each emphasizing two to three criteria extracted from the fundamental task characteristics: action-execution correctness, preservation of non-target regions, and overall realism for Action; positional accuracy, occlusion consistency, and global harmony for Location; and plausibility of viewpoint transformation, geometric coherence, and detail preservation for Viewpoint. All prompts instruct GPT-4o to return a standardized JSON object of the form {"score": <0-10>, "reason": "<...>"}. Subsequently, each image pair is processed by GPT-4o precisely once with the decoding temperature fixed at 0, yielding deterministic and consequently fully reproducible results.
Appendix E Human Annotations
Annotation Details. To complete our benchmark development, we engaged four expert annotators and one quality assurance specialist to supervise all human verification processes throughout the benchmark construction. This comprehensive oversight encompassed filtering the initial datasets, validating generated images, and most critically, formulating precise instructions. During the instruction annotation phase, annotators were provided with the original images, edited images, and corresponding masks, and were tasked with crafting instructions based on this multimodal information. All annotators underwent extensive training, including exposure to exemplary cases, and received iterative feedback to progressively enhance their performance. To ensure annotation fidelity and inter-annotator consistency, the quality assurance specialist conducted systematic evaluations of all outputs. Significantly, all procedures were implemented in a double-blind framework to mitigate potential experimental biases.
Participant Disclosure and Consent. Annotators were comprehensively briefed regarding the purpose of the annotation task, their unconditional right to withdraw participation at any juncture, and the exclusive research-oriented utilization of their annotations. No personally identifiable information was collected during this process. All annotators provided documented informed consent prior to their participation in the study.
Compensation. Annotators received remuneration at a rate commensurate with or exceeding the local minimum wage standards.
Appendix F Statement of Limitation, Ethical Concern and Broader Impact
Limitation. A significant constraint of our methodology resides in the substantial computational overhead and procedural complexity of the integrated pipeline. The framework encompasses multiple sequential stages of model-based processing followed by comprehensive human evaluation, necessitating considerable computational resources and expert human intervention. This intricate design imposes reproducibility challenges, hampers scalability, and restricts the exhaustive coverage of the complete spectrum of compositional editing tasks. Furthermore, components leveraging large language models are inherently bounded by contemporary model limitations, potentially compromising performance on tasks demanding sophisticated reasoning or precise visual-linguistic integration.
Ethical Consideration. While our instruction-based image editing framework demonstrates robust capabilities, it elicits ethical considerations regarding potential misappropriation and privacy implications. The capacity to generate photorealistic manipulations may be exploited to create deceptive content, potentially exacerbating misinformation dissemination if inadequately regulated. Moreover, the utilization of real-world imagery introduces non-negligible risks of inadvertently exposing personally identifiable information, despite rigorous anonymization protocols. Ensuring responsible deployment necessitates implementing comprehensive safeguards, including content verification mechanisms, stringent data protection frameworks, and transparent operational guidelines.
Broader Impact. CompBench establishes a rigorous evaluation framework for instruction-guided image editing, facilitating systematic assessment of complex editing capabilities in multimodal large language models. This contribution will accelerate the development of reasoning-aware and controllable editing systems, enhance model performance across visual understanding and generation domains, and expand the practical applicability of large-scale models in diverse real-world contexts, including creative design processes, digital content production, and interactive artificial intelligence assistants.